# Coach not crutch: AI assistance can enhance rather than hinder skill development

### v3

@@ Abstract
Does using AI make you stupid? Contrary to predictions by lay forecasters, participants randomly assigned to practice writing with access to an AI tool improved more on a writing test one day later compared to writers assigned to practice without AI—despite exerting less effort. Likewise, whereas forecasters predicted that AI would be less helpful, and less worth paying for, than personalized feedback from experienced human editors, we found the converse: practicing with AI increased skill more than feedback from human editors (or simply googling examples and tips). The results of a third experiment suggest that AI may teach by example: writers merely shown an AI-generated cover letter (without the opportunity to practice) performed as well as writers who practiced writing with the original AI tool. Collectively, the findings of these pre-registered studies constitute an existence proof that by providing personalized, high-quality examples, AI can improve, rather than undermine, learning.
@@
[…]
@@
Generative AI (henceforth AI) tools are increasingly powerful and prevalent (1), and there is mounting evidence that they can dramatically boost performance. For example, working side-by-side with AI as a copilot has been shown to increase both quality and speed in a variety of professional writing tasks (e.g., emails, memos, short reports) (2, 3, 4).
Nevertheless, there is growing concern that AI tools will be used as a crutch, providing immediate gains in performance at the expense of long-run development of human capital (5, 6, 7). For instance, in a 2024 poll, 62% of surveyed adults predicted that Generative AI would “lead to humans becoming less intelligent” (8). In January 2023, New York City public schools banned ChatGPT, citing “concerns about negative impacts on student learning (9).” When this ban was lifted three months later, it was not because of AI’s potential to scaffold learning, but instead because of the “reality that students are participating in and will work in a world where understanding Generative AI is crucial (10).” However, teachers remain worried: in a nationally representative poll in May 2024, four times as many K-12 educators judged the use of AI tools as net harmful (24%) than net beneficial (6%) (11).
Concerns that using AI tools hinders learning (while increasing short-term performance) are justified for at least three reasons. First, AI systems based on large language models like GPT-4 have been shown to confidently assert erroneous facts (i.e., hallucinations (12)), make reasoning and arithmetic errors (13), and complete other tasks with varying degrees of accuracy.
Second, the fluent and instantaneous solutions AI tools generate may contribute to an illusion of mastery. To the extent users conflate the skills of an AI tool with their own, they may be less likely to seek feedback and improve. Prior research has found that searching for information on the Internet, for example, creates an illusion whereby people conflate knowledge outside their heads with what they personally know themselves (14).
Third, technological tools reduce the learner’s need to be cognitively engaged with the task at hand. For instance, knowing that one can search for a fact on a computer has been shown to reduce memory for that fact, instead encouraging recall of how to search for it (15). And drivers who use GPS tend to have worse hippocampal-dependent spatial memory, both cross-sectionally and longitudinally (16, 17). When AI-powered tools instantaneously produce turn-key solutions for complex cognitive tasks, they may be especially detrimental to learning. It is, after all, tempting to copy and paste the output of an AI tool without even laying eyes on it.
On the other hand, using AI tools to assist us in our work may, in fact, help us develop our own skills. In particular, the current generation of AI tools may teach by example, offering high-quality and personally tailored demonstrations of abstract principles that are otherwise difficult to grasp. Classic research shows that worked examples of math problems (i.e., not just answers but the step-by-step process by which problems are solved) scaffold learning more effectively than explanations alone (18, 19). It seems likely that, compared to textbooks, conventional computer tutoring programs (e.g., (20)), and human teachers, AI tools may provide higher quality, just-in-time examples exquisitely tailored to a learner’s specific needs. Indeed, it’s possible that in both quality and didactic utility, AI examples (e.g., of an excellent cover letter) might surpass even those provided by domain experts (e.g., professional writers and editors). Thus, if the upside of exposure to excellent examples tailored to the learners’ needs outweighs the downside of diminished engagement, AI tools may support skill development.
The scant research on how AI tools influence skill development has yielded mixed results. Some studies report that interacting with AI tools improves skill on subsequent tests in which AI tools are not available (21, 22), while others have shown null or even negative effects (23, 24, 22). Notably, these studies examine AI tutors, chatbots, or explanations explicitly designed to support learning, rather than simply providing solutions, as is typical in real-world use. Further, they focused exclusively on mathematics and computer programming.
In this investigation, we ask whether AI tools can support skill development, above and beyond merely improving performance while using the AI tool. We focus on writing—the most common use of AI at work, as ranked in a nationally representative survey of American adults in August 2024 (26). In our experimental paradigm all participants are given a baseline writing test (i.e., revising a poorly written cover letter), followed by a short lesson on five evidence-based strategies for professional writing (25, 27). Next, participants are randomly assigned to different types of practice (e.g., practice with access to an AI tool, practice without access to an AI tool, practice with feedback from professional editors, and practice with Google Search to find examples and tips). Finally, all participants complete an incentivized test in which they rewrite cover letters without access to AI.
@@
In Study 1, forecasters presented with this design were twice as likely to predict that practicing writing with the assistance of AI would impair learning compared to practicing without it. In Study 2, however, participants who practiced with AI actually learned more (i.e., wrote better cover letters during the test phase) than participants who practiced without AI—an advantage that persisted in a one-day follow-up test. In Study 3, forecasters predicted that feedback from AI would be less helpful than feedback from professional editors and, accordingly, were less willing to pay for it. However, in Study 4, participants who practiced with AI learned more than participants who practiced with feedback from professional editors (or those who practiced with Google Search to find examples and tips). Finally, in Study 5, we explored the mechanism for these learning gains by introducing an example-only condition. Participants who saw an AI-generated example (but did not have an opportunity to practice) improved their writing as much as participants who practiced with AI; benefits again persisted in a one-day follow-up test.
Study 1: Forecasting the effect of AI on learning
We showed N = 150 participants screenshots of a random assignment study with three conditions. We asked them to rank-order these conditions according to how much they predicted future participants would learn in each.
Lay forecasters predicted that practicing with AI would hinder learning
Confirming our pre-registered hypothesis, nearly twice as many forecasters (64.7%) ranked practicing alone above practicing writing with access to an AI tool as the converse (35.3%, (1) = 12.9, p < .001), see Figure 3. Participants made this prediction regardless of self-reported experience with AI (OR = 0.72, p = .096) or any other measured demographic characteristic (ps > .05). See Table S2
In open-ended responses, forecasters who were pessimistic about the effect of the AI tool on learning speculated that it would crowd out effort (e.g., “Practicing alone would force more recall and problem-solving skills, while AI essentially just gives them the answer.”, “I think oftentimes using AI impedes the learning process because it’s the ‘easy way.”’). Those with positive views, on the other hand, cited the possibility of AI providing insights or examples that would be otherwise unavailable (“As much as I hate AI, I do not believe you can improve in any manner if you do not have examples or other ways of learning, and AI can provide this.”)
@@
Study 2 tested whether the predictions of Study 1 forecasters were accurate. Specifically, N = 2,238 participants completed a baseline questionnaire and pretest (rewriting a poorly written cover letter), followed by a lesson introducing five principles of effective writing (i.e., Less is more, Make reading easy, Design for easy navigation, Use enough formatting but no more, Make responding easy) (25). Next, participants were randomly assigned to one of three practice conditions: (1) rewriting a new cover letter with an AI writing tool that instantly revised text based on these principles, (2) rewriting the new cover letter without the AI tool, or (3) a no-practice control. At the end of the session, all participants completed a test of writing skill (rewriting a new cover letter without access to the AI writing tool) and an exit questionnaire. One day later, all participants were invited to complete a similar incentivized test of writing skill. See Figure 2.
We assessed writing skill in two ways. First, we used GPT-4o to rate cover letters on each of the principles of effective writing and then averaged these ratings to produce a summary score (Cronbach’s = .81). This operationalization demonstrated convergent validity with ratings by human research assistants who rated a random subsample of n = 100 cover letters on the same principles (r = .70, p < .001). Second, we recruited participants naive to this investigation to read two letters and select the one more likely to secure its writer a job interview. These two operationalizations converged: preferred cover letters obtained higher AI ratings of writing skill (rs = .29).
@@
Consistent with other studies demonstrating the productivity benefits of AI tools (3, 2), participants given access to the AI writing tool produced cover letters during the practice phase that were dramatically higher in quality than participants without access (d = 1.01, p <.001).
The learning advantage of having practiced with AI was evident in the test phase: Consistent with our preregistered hypothesis, participants who had practiced with the AI tool produced higher-quality writing than did participants who either had practiced without the AI tool (d = .38, p <.001) or who had not practiced at all (d = .47, p <.001). See Figure S3. Likewise, cover letters written by participants who had practiced with AI were more likely to secure a hypothetical job interview than cover letters by participants who had practiced without AI (.54 vs. .50, p = .002) or had not practiced at all (.54 vs. .47, p < .001). See Figure 4B.
@@
Compared to participants who practiced alone, participants who practiced with the AI tool spent 0.44 fewer minutes during the practice phase (3.73 vs. 4.17; d = .12, p = .025, quantity of effort), logged roughly a quarter as many total keystrokes (26% d = .44, p <.001) and logged 65% fewer keystrokes per minute (d = .22, p <.001, intensity or quality of effort), and self-reported expending less effort during practice (d = .31, p <.001).
Nevertheless, it would be inaccurate to label writers practicing with AI as entirely disengaged. Rather than copying and pasting the AI tool’s output, the majority of participants practicing with AI chose to interact with the task for over 3 minutes, and 95% made at least one edit to the AI-generated output before final submission. See Figure S4 in Supplementary Information for details.
@@
In the exit questionnaire, participants who had practiced with AI reported learning about as much as participants who had practiced without it or not practiced. Likewise, they judged their writing skill similarly to those who had either practiced alone or done no practice at all (|ds| 0.10, ps >.05). When asked if they wanted to see additional feedback, they were about as likely to say yes as participants who had practiced without AI (.64 vs. .60, p = .167), and slightly less likely than those who had not practice at all (.65 vs. .60, p = .039). See Section C4 in the Supplementary Information for details.
The benefits of practicing with AI were just as large a day later
@@
Study 3: Willingness to pay for AI vs. human editors
Given unwarranted pessimism about the pedagogical power of AI, in Study 3 we examined whether people underinvest in AI as a learning resource. Specifically, we asked N = 150 people to imagine they were participants in Study 2 and asked them how much they would be willing to pay for feedback from the AI tool vs. feedback from experienced human editors.
Participants were willing to pay more for feedback from expert human editors than for AI feedback
Out of a maximum of a dollar (the total amount of the hypothetical bonus participants could earn), participants on average were willing to pay 28 cents for the AI tool vs. 38 cents for feedback from an experienced human editor (t(149) = 4.83, p <.001, d = 0.39).
Participants predicted that editorial feedback would be more beneficial
Participants predicted that experienced human editors (M = 6.68 on a 0—10 Likert scale) would be more effective in helping them improve their writing, compared to feedback from an AI tool (M = 5.64; t(148) = 4.77, p <.001, d = 0.39).
Predicted effectiveness mediated underinvestment in AI feedback
People’s beliefs about how effective AI and human editor feedback would be predicted willingness to pay (b = 0.06, SE = 0.01, p < .001) and accounted for most of the preference for human editors (indirect effect = 0.06, p < .001). Once these predictions were accounted for, the remaining preference for human feedback was reduced by 60% (c’ = 0.04, SE = 0.02, p = .054). See Figure 5 and Section D1 of Supplementary Information for details.
Study 4: Practicing with AI vs. with expert feedback (vs. with Google Search)
Participants (N = 2,997) completed a baseline questionnaire and pretest, followed by the same lesson on effective writing as in Study 2, and then practiced rewriting a new cover letter on their own. Next, participants were randomly assigned to one of three conditions: (1) practice with AI, (2) practice with feedback from professional human editors, (3) practice with Google Search. Letters written by participants assigned to the second condition were given to professional editors averaging 25 years of professional editorial and writing experience. These 49 editors, recruited in partnership with the Journalists’ Resource at Harvard Kennedy School, provided personalized revisions and constructive feedback. The following day, all participants attempted to improve the cover letter they had previously submitted, using the resources afforded by their assigned condition. Finally, all participants completed a writing test (rewriting a cover letter without access to AI) and an exit questionnaire.
@@
During the test phase, in which participants were asked to write a new cover letter without access to any outside resources, participants who had practiced with the AI tool produced higher-quality cover letters than those who had practiced with Google Search (d = 0.46, p <.001), and even outperformed those who had practiced with personalized feedback from professional editors (d = 0.20, p <.001). See Figure 6A.
@@
A possible concern is that reviewing many letters consecutively caused editors to get tired and give worse feedback over time. This could suggest that our experiment compared AI against a “strawman” version of human expertise. However, analysis of the editors’ performance revealed the opposite pattern: while editors became more efficient, spending less time per letter as the task progressed (b = 0.387, p < .001), the quality of their edits and feedback actually improved (b = 0.013, p < 0.001). Thus, the AI tool’s advantage over professional editors cannot be attributed to a decline in editorial performance. See Section E2 of Supplementary Information for details.
@@
During the practice phase, participants with access to AI spent about the same time as those who practiced with editor feedback (difference = 0.43 minutes, d = 0.09, p > .05), but about 1.18 fewer minutes than those who practiced with Google Search (d = 0.32, p < .001). Nevertheless, participants across all three groups reported similar levels of subjective effort (ps > .05). Likewise, participants who practiced with AI logged about as many keystroles as those who practiced with editor feedback (68% fewer, d = 0.08, p > .05) and 76% fewer keystrokes than those who practiced with Google Search (d = 0.32, p < .001). They also typed at a slower rate than participants who practiced with editor feedback (d = 0.41, p < .001) or those who practiced with Google Search (d = 0.34, p < .001).
@@
Participants who had practiced with AI did not report learning more than those who had practiced with editor feedback (d = 0.07, p > .05), nor did they consider their writing skills to be better (d = 0.02, p > .05). Both groups, however, correctly recognized that they had learned more than those who had practiced with Google Search (ds = .28 and .20, respectively), and evaluated their writing skills accordingly (ds = .11 and .12, respectively). Participants who had practiced with AI and those who had practiced with editor feedback requested feedback on their test submissions at similar rates (71% and 70%, respectively). Both groups were more likely to request feedback than those who had practiced with Google search (64%, p = .003 and .007, respectively). See Section E6 in the Supplementary Information for details.
@@
Study 5: Examples as the mechanism for AI learning gains
To better understand what drives the benefit of AI practice, in Study 5 (N = 2,003), we preregistered a replication and extension in which we replaced the no-practice condition of Study 2 with an example-only condition. In this condition, we showed participants an AI-generated writing example that they could not edit. To the extent that the benefit of practicing with AI was driven by exposure to examples, the example-only condition should improve performance in the test phase as much as the practice with AI condition.
@@
Participants who had merely seen an AI-generated example also improved more in writing skill than those who had practiced without AI (d = .37, p < .001), and produced letters that were relatively more likely to secure them interviews (.52 vs. .47, p = .007). Notably, they improved as much as participants who had practiced with the AI tool (and could edit its output, d = .03, p = .883), and were offered hypothetical interviews at similar rates as them (.51 vs. .52, p = .561). See Figures S10 and 7B.
@@
As in Studies 2 and 4, despite learning more, participants who had practiced with AI or had merely seen an AI-generated example reported learning similar amounts to those who practiced without AI (ps > .05) and rated their writing skill after practice at comparable levels (ps > .05). Moreover, all participants requested feedback at similar rates (proportions ranged from 63% to 67%). See Section F4 in the Supplementary Information for details.
@@
When we recontacted a subsample of participants (n = 800) one day later, the majority responded (n = 633, 80%); the attrition rates ranged from 17% to 24% and did not differ by condition ( (2) = 4.56, p = .102). The effect remained robust after 24 hours. Participants who had practiced with the AI-tool (d = .29, p = .006) and participants who had simply seen an AI example (d = .32, p = .003), both continued to outperform those who had practiced without the tool. Participants who had merely seen an AI example performed as well as those who had practiced with AI (d = .02, p = .830). As in Study 2, the magnitude of the follow-up effect size was not significantly different from the immediate test phase effect size (condition time interactions were non-significant (ps > .96). See Figure 7A and Section F5 of Supplementary Information for details.

### v4

# Coach not crutch: Evidence that AI can improve writing skill despite reducing effort

@@ Abstract
In a series of highly-powered empirical studies, we examine the intuition that by sparing effort, using AI inevitably hinders learning. First, in a nationally representative survey of young adults, the majority expressed the view that using AI makes people lazier and less capable. Next, in a random-assignment experiment, we gave participants a tutorial on best practices in professional writing, then provided one group with access to an AI writing tool and asked another to practice writing on their own. Those who practiced with AI indeed exerted less effort while practicing—yet wrote better cover letters in no-AI writing tests. In a second experiment with more rigorous control conditions, access to AI improved writing more than either googling cover letter examples and tips or receiving personalized feedback on their practice letters from experienced human editors. A third experiment explained these learning gains by showing that AI can teach by example: participants who viewed a cover letter that had been revised by the AI tool but did no further practice improved their writing as much as those who practiced writing with the original AI tool. Collectively, these pre-registered experiments suggest that AI can exert opposing effects on effort and learning rate—making it possible in at least some cases to work less and learn more.
@@
Human Capital Development | Generative AI | Skill AcquisitionThis manuscript was compiled on February 18, 2026
@@
Generative AI (henceforth AI) tools are increasingly powerful and prevalent (bubeck2023), and there is mounting evidence that they can dramatically boost performance. For example, working side-by-side with AI as a copilot has been shown to increase both quality and speed in a variety of professional writing tasks (e.g., emails, memos, short reports) (noy2023, dellacqua2023, wiles2023algorithmic).
Nevertheless, advances in AI in recent years have been accompanied by concerns that immediate gains in performance are offset by the subsequent erosion of human capital (hofman2023sports, puntoni2021, borges2024). For example, in a 2024 poll of American adults, 62% predicted that generative AI would “lead to humans becoming less intelligent” (hawkins2024between). In a 2024 survey of K–12 educators, four times as many judged AI to be net harmful (24%) as net beneficial (6%) (lin2024a).
Concerns that AI hinders learning are justified for at least three reasons. First, AI sometimes confidently asserts erroneous facts (i.e., hallucinations (ji2023survey, yuan2023well)). Second, using AI may create an illusion of mastery (fisher2015), particularly if users conflate its skilled responses with their own. Third and most important, using AI may reduce human effort because the more AI solves your problems, the less you have to think about them yourself (kosmyna2025your, bastani2024). All things being equal, less effort results in less learning (bjork2011making, james1983talks).
Analogous deskilling effects have been documented with prior technologies. In cross-sectional and longitudinal studies, drivers who rely on GPS exhibit poorer hippocampal-dependent spatial memory (maguire2000, griesbauer2022london). Likewise, knowing that information can be easily retrieved from the internet reduces the likelihood that one will commit it to memory. (sparrow2011). AI may be uniquely harmful because, unlike search engines or GPS, it can replace the process of thinking itself—not just the storage or retrieval of facts.
The assumption that diminished effort inevitably reduces learning is shaping the design of real-world AI learning tools. For example, when Khan Academy released its AI tutor Khanmigo, it announced on its website that it would prevent students from seeing full solutions: “Khanmigo challenges you to think critically and solve problems without giving you direct answers111emphasis added” (khanmigo2025).
Skill acquisition, however, depends on more than effort; learning rate also matters (duckworth2015mechanics). A wide array of factors have been shown to influence how quickly students learn in conventional educational settings (wisniewski_power_2020, bloom_2_1984, sana_interleaving_2022). One feature of optimal learning environments with special relevance to the affordances of generative AI is the availability of specific examples that illustrate abstract principles that are otherwise difficult for the learner to grasp and apply. Separate from its influence on effort, therefore, AI might change the quality of the learning environment. See Figure 1. The net effect of effort and learning environment on skill acquisition remains an open empirical question.
Here, we focus on writing—the most common use of AI at work. In a nationally representative sample of American adults in 2024, writing tasks were the most frequent workplace application (bick2024rapid). This pattern is corroborated by large-scale usage data: In an analysis of over 1 million ChatGPT conversations writing constituted 24% of all messages and 40% of work-related messages, making it the dominant occupational use case (chatterji_how_nodate). Notably, approximately two-thirds of writing-related queries involved modifying existing text rather than generating novel content.
We know of only one study (kosmyna2025your) that has experimentally manipulated access to AI (vs. no access) during a writing task. Fifty-four undergraduates at four selective universities were randomly assigned to condition and then asked to write essays: 18 participants wrote with access to AI, 18 wrote without access to AI, and 18 wrote with access to Google Search. Although a no-AI test of writing skill was not administered, compared to the other two conditions, participants given access to AI were less likely to recall a verbatim sentence from their essays immediately after writing them.
We developed an experimental paradigm to determine whether using AI as a productivity tool for writing cover letters enhances or erodes subsequent writing skill. To begin, all participants complete a baseline questionnaire and writing pretest (i.e., revising a poorly written cover letter). Next, as shown in Figure 4, all participants are taught five evidence-based principles for professional writing (i.e., Less is more, Make reading easy, Design for easy navigation, Use enough formatting but no more, Make responding easy) (rogers2023, shulman2024) and are shown a domain-general example for each principle. After this lesson, participants are randomly assigned to different types of practice (e.g., practice with access to an AI tool, practice without access to an AI tool, practice with feedback from professional editors, and practice with Google Search to find examples and tips). Finally, all participants rewrite a new cover letter without access to AI (and with copy-pasting functionality disabled). Each participant’s submitted cover letter is rated on adherence to each of the five writing principles, respectively. To gauge the ecological validity of the resulting composite index, we recruited a separate sample of raters (blind to condition) who read randomly chosen pairs of cover letters from the main experiment and were asked to “choose the cover letter that would make you more likely to offer an interview to the candidate.” See Methods for details.
@@
In this investigation, we examine contemporary assumptions about learning and AI and empirically test the accuracy of these beliefs. In Study 1, a nationally representative survey of young adults in the U.S. in October 2025, the majority endorsed the view that using AI makes people lazier, less smart, and less educated. In Study 2, participants who practiced writing cover letters with AI actually learned more than participants who practiced without AI—an advantage that persisted in a one-day follow-up test. A separate sample of lay forecasters predicted the opposite result. In Study 3, again against forecaster predictions, participants who practiced with AI learned more than participants who practiced with feedback from professional editors (or those who practiced with Google Search to find examples and tips). Finally, in Study 4, we explored the mechanism for these learning gains by introducing an example-only condition. Participants who viewed an AI-generated example (but did not have an opportunity to edit it) improved their writing as much as participants who practiced with AI; benefits again persisted in a one-day follow-up test.
Study 1: Public perceptions of AI and learning
In partnership with the Gallup Organization, we asked a nationally representative sample of American adults aged 18 to 28; (N = 2,637) in October 2025 about AI use in the past month and attitudes about AI in general.
Consistent with 2024 polls, writing is the most common use case for AI (45% of all adults surveyed) after using AI as a replacement for Google (64%) and ahead of generating images (35%) or help with math (23%). In general, AI is used more commonly for work (53%) than for personal reasons like life advice (28%) or companionship (17%).
Participants expressed concern that “AI makes people lazier” (79%, One sample d = 0.66, p < .001) and, to a more moderate degree, that “AI makes people learn less” (62%, One sample d = 0.34, p < .001) and “AI makes people less smart” (63%, One sample d = 0.30, p < .001). See Figure 5. See Section B in the Supplementary Information for details.
@@
In Study 2 (N = 2,238), we preregistered an experiment in which participants completed a baseline questionnaire and pretest, followed by a lesson on effective writing. They were then randomly assigned to one of three practice conditions: (1) rewriting a new cover letter with an AI writing tool that instantly revised text based on these principles, (2) rewriting the new cover letter without the AI tool, or (3) a no-practice control. Because forgetting occurs rapidly—with substantial decay even within 24 hours (e.g., (roediger_test-enhanced_2006, murre2015replication)), we invited all participants to complete a similar incentivized test of writing skill one day later. See Figure 3.
@@
Consistent with our preregistered hypothesis, participants who had practiced with the AI tool wrote better cover letters than participants who either had practiced without the AI tool (d = .38, p <.001) or who had not practiced at all (d = .47, p <.001). See Figure S3. Likewise, cover letters written by participants who had practiced with AI were more likely to secure a hypothetical job interview than cover letters by participants who had practiced without AI (.54 vs. .50, p = .002) or had not practiced at all (.54 vs. .47, p < .001). See Figure 6B.
@@
Compared to participants who practiced alone, participants who practiced with the AI tool spent 0.44 fewer minutes during the practice phase (3.73 vs. 4.17; d = .12, p = .025), logged roughly a quarter as many total keystrokes (26% d = .44, p <.001) and logged 65% fewer keystrokes per minute (d = .22, p <.001), and self-reported expending less effort during practice (d = .31, p <.001).
Nevertheless, it would be inaccurate to label writers practicing with AI as entirely disengaged. Rather than simply copying and pasting the AI tool’s output, the majority of participants practicing with AI chose to interact with the task for over 3 minutes, and 95% made at least one edit to the AI-generated output before final submission. See Figure S4 in Supplementary Information for details.
@@
In the exit questionnaire, participants who had practiced with AI reported learning about as much as participants who had practiced without it or not practiced. Likewise, they judged their writing skill similarly to those who had either practiced alone or done no practice at all (|ds| 0.10, ps >.05). When asked if they wanted to see additional feedback, they were about as likely to say yes as participants who had practiced without AI (.64 vs. .60, p = .167), and slightly less likely than those who had not practiced at all (.65 vs. .60, p = .039). See Section C4 in the Supplementary Information for details.
The benefits of practicing with AI were just as large one day later
@@
Forecasters predicted AI would reduce effort and harm learning
We showed N = 150 participants screenshots of this design and asked them to rank-order the study conditions according to how much they predicted participants would learn in each. Confirming our pre-registered hypothesis, nearly twice as many forecasters (64.7%) ranked practicing alone above practicing writing with access to an AI tool as the converse (35.3%, (1) = 12.9, p < .001), see Figure S6. Participants made this prediction regardless of self-reported experience with AI (OR = 0.72, p = .096) or any other measured demographic characteristic (ps > .05). See Table S16
Forecasters who were pessimistic about the effect of the AI tool on learning speculated that it would crowd out effort (e.g., “Practicing alone would force more recall and problem-solving skills, while AI essentially just gives them the answer.”, “I think oftentimes using AI impedes the learning process because it’s the ‘easy way.”’). Those with positive views, on the other hand, cited the possibility of AI providing insights or examples that would be otherwise unavailable (“As much as I hate AI, I do not believe you can improve in any manner if you do not have examples or other ways of learning, and AI can provide this.”)
Study 3: Practicing with AI vs. with expert feedback (vs. with Google Search)
In Study 3 (N = 2,997), we preregistered an experiment testing whether AI was more helpful than two ecologically valid comparison conditions. Participants (N = 2,997) completed a baseline questionnaire and pretest, followed by the lesson on effective writing, and then practiced rewriting a new cover letter on their own. Next, participants were randomly assigned to one of three conditions: (1) practice with AI, (2) practice with feedback from professional human editors, (3) practice with Google Search. Letters written by participants assigned to the second condition were given to professional editors averaging 25 years of professional editorial and writing experience. These 49 editors, recruited in partnership with the Journalists’ Resource at Harvard Kennedy School, provided personalized revisions and constructive feedback. The following day, all participants attempted to improve the cover letter they had previously submitted, using the resources afforded by their assigned condition. Finally, all participants completed a writing test (rewriting a cover letter without access to AI) and an exit questionnaire.
@@
During the test phase, in which participants were asked to write a new cover letter without access to any outside resources, participants who had practiced with the AI tool wrote higher-quality cover letters than those who had practiced with Google Search (d = 0.46, p <.001), and even outperformed those who had practiced with personalized feedback from professional editors (d = 0.20, p <.001). See Figure 7A.
@@
A possible concern is that reviewing many letters consecutively caused editors to get tired and give worse feedback over time. This could suggest that our experiment compared AI against a “strawman” version of human expertise. However, analysis of the editors’ performance revealed the opposite pattern: while editors became more efficient, spending less time per letter as the task progressed (b = 0.387, p < .001), the quality of their edits and feedback actually improved (b = 0.013, p < 0.001). Thus, the AI tool’s advantage over professional editors cannot be attributed to a decline in editorial performance. See Section D2 of Supplementary Information for details.
The quality of edits provided by AI was higher than that of edits from professional human editors. We used the AI tool to rewrite each of the texts assigned to the editor pool, creating a counterfactual of how AI would have edited those letters. Relative to the counterfactual, letters rewritten with the AI tool were rated higher for writing quality (d = 1.66, p < .001), and when shown to independent raters, they were considerably more likely to secure a hypothetical job interview (d = 0.75, p < .001). Thus, AI users encountered examples that more closely adhered to the writing principles than those provided by human editors, helping explain the learning advantage associated with AI use.
@@
During the practice phase, participants with access to AI spent about the same time as those who practiced with editor feedback (difference = 0.43 minutes, d = 0.09, p > .05), but about 1.18 fewer minutes than those who practiced with Google Search (d = 0.32, p < .001). Nevertheless, participants across all three groups reported similar levels of subjective effort (ps > .05). Likewise, participants who practiced with AI logged about as many keystrokes as those who practiced with editor feedback (68% fewer, d = 0.08, p > .05) and 76% fewer keystrokes than those who practiced with Google Search (d = 0.32, p < .001). They also typed at a slower rate than participants who practiced with editor feedback (d = 0.41, p < .001) or those who practiced with Google Search (d = 0.34, p < .001).
@@
Participants who had practiced with AI did not report learning more than those who had practiced with editor feedback (d = 0.07, p > .05), nor did they consider their writing skills to be better (d = 0.02, p > .05). Both groups, however, correctly recognized that they had learned more than those who had practiced with Google Search (ds = .28 and .20, respectively), and evaluated their writing skills accordingly (ds = .11 and .12, respectively). Participants who had practiced with AI and those who had practiced with editor feedback requested feedback on their test submissions at similar rates (71% and 70%, respectively). Both groups were more likely to request feedback than those who had practiced with Google search (64%, p = .003 and .007, respectively). See Section D6 in the Supplementary Information for details.
@@
Forecasters predicted AI would be less helpful than human editors and were less willing to pay for it
We asked N = 150 people to imagine they were participants in our experiment and asked them how much they would be willing to pay for feedback from the AI tool vs. feedback from experienced human editors. Out of a maximum of a dollar (the total amount of the hypothetical bonus participants could earn), participants on average were willing to pay 28 cents for the AI tool vs. 38 cents for feedback from an experienced human editor (t(149) = 4.83, p < .001, d = 0.39). Participants predicted that experienced human editors (M = 6.68 on a 0—10 Likert scale) would be more effective in helping them improve their writing, compared to feedback from an AI tool (M = 5.64; t(148) = 4.77, p < .001, d = 0.39). People’s beliefs about the effectiveness of AI and human editor feedback predicted willingness to pay (b = 0.06, SE = 0.01, p < .001) and accounted for most of the preference for human editors (indirect effect = 0.06, p < .001). Once these predictions were accounted for, the remaining preference for human feedback was reduced by 60% (c’ = 0.04, SE = 0.02, p = .054). See Figure S11 and Section D9 of Supplementary Information for details.
Study 4: Examples as the mechanism for AI learning gains
To better understand how using AI improves writing skill, in Study 4 (N = 2,003), we preregistered a replication and extension in which we replaced the no-practice condition of Study 2 with an example-only condition. In this condition, we showed participants an AI-generated writing example that they could not edit. To the extent that the benefit of practicing with AI was driven by exposure to examples, the example-only condition should improve performance in the test phase as much as the practice with AI condition.
@@
Participants who had merely seen an AI-generated example also improved more in writing skill than those who had practiced without AI (d = .37, p < .001), and produced letters that were relatively more likely to secure them interviews (.52 vs. .47, p = .007). Notably, they improved as much as participants who had practiced with the AI tool (and could edit its output, d = .03, p = .883), and were offered hypothetical interviews at similar rates as them (.51 vs. .52, p = .561). See Figures S12 and 8B.
@@
As in Studies 2 and 4, despite learning more, participants who had practiced with AI or had merely seen an AI-generated example reported learning similar amounts to those who practiced without AI (ps > .05) and rated their writing skill after practice at comparable levels (ps > .05). Moreover, all participants requested feedback at similar rates (proportions ranged from 63% to 67%). See Section E4 in the Supplementary Information for details.
@@
When we recontacted a subsample of participants (n = 800) one day later, the majority responded (n = 633, 80%); the attrition rates ranged from 17% to 24% and did not differ by condition ( (2) = 4.56, p = .102). The effect remained robust after 24 hours. Participants who had practiced with the AI tool (d = .29, p = .006) and participants who had simply seen an AI example (d = .32, p = .003), both continued to outperform those who had practiced without the tool. Participants who had merely seen an AI example performed as well as those who had practiced with AI (d = .02, p = .830). As in Study 2, the magnitude of the follow-up effect size was not significantly different from the immediate test phase effect size (condition time interactions were non-significant (ps > .96). See Figure 8A and Section E5 of Supplementary Information for details.